# A linear-time algorithm computing the resident fitness in interacting trajectories

### v1

@@ 4.3 Proof of Proposition 12
Since $\Lambda(\alpha):=\log\mathbb{E}(\mathrm{e}^{\alpha E_{1}})<\infty$ for $\alpha>0$ small (which also follows from the stochastic domination by a geometric random variable), Cramér’s theorem [8, Section 2.2] implies that if $\delta^{\prime}$ is sufficiently small, then
@@
and together with the fact that $\Lambda^{\prime}(0)=m_{1}$ and $\Lambda^{*}(m_{1})=0$, the strict negativity in (4.3) follows from the assertion of [8, Exercise 2.2.24] for $\delta^{\prime}>0$ small enough. We conclude that for all sufficiently large $t$,
@@
as wanted.  $\Box$
@@ 5.1 Ancestral relations and fixations
According to [13, Section 6], for $i\geq 0$ we say that $h_{i}$ fixes (or reaches fixation) if there exists $t_{0}>0$ such that for all $t>t_{0}$, the trajectory $h_{i}$ is an ancestor (i.e., parent, or parent of parent etc.) of all trajectories $h_{j}$ such that $h_{j}(t)>0$. If the $i$-th trajectory is alone in the system with its parent, $h_{i}$ reaches height 1 at time $t_{i}+\frac{1}{a_{i}}$ and immediately fixes, but due to clonal interference with other trajectories, this is not always the case: The slope of the $i$-th trajectory may change due to resident changes after time $t_{i}$.
@@
Given our algorithm in the proof of Theorem 11 for the resident fitness, for finite $n$ we can now determine in linear time which trajectories fix. Indeed, the trajectories that fix are precisely the ultimate resident and its ancestors, and this ancestral line can be found in $O(n)$ time using linear search. We can search for the parent of the ultimate resident starting from the end of the list in the output of the algorithm, and then we can search for the parent of this parent continuing backwards etc.
5.2 Computing height functions and values of the resident fitness
@@
Given the form of the output of the algorithm in Section 4.2, which we can convert from a doubly linked list into an array at the end of algorithm within $O(n)$ time, for any fixed time $t>0$ we can determine the resident fitness and resident type at time $t$ within a runtime of $O(\log n)$ using binary search. Next, if we want to compute $h_{i}(t)$ for some $i$ (where $t\geq t_{i}$), we can first compute the resident fitness values $f(t_{i})$ and $f(t)$ and the resident types $\varrho(t_{i})$ and $\varrho(t)$ at times $t_{i}$ and $t$. Then if say $s_{j}<t_{i}<s_{j+1}$ and $s_{k}<t<s_{k+1}$, we know that in the continued lines representation, the $i$-th ray $s\mapsto l_{i}(s)$ starts with slope $f(t_{i})+a_{i}$ at height $y_{j}+(t_{i}-s_{j})b_{m_{j}}-1$ at time $t_{i}$. Now, in the original system of interacting trajectories, we have $h_{i}(t)=l_{i}(t)-(l_{\varrho(t)}-1)=l_{i}(t)-(l_{m_{k}}(t)-1)$ if this quantity is positive and $h_{i}(t)=0$ otherwise. Thus, $h_{i}(t)$ can also be computed in $O(\log n)$ time (for fixed $t$). Moreover, the entire piecewise linear trajectory $t\mapsto h_{i}(t)$ can be determined in $O(n)$ time as follows: at time $t_{i}$ it starts with slope $a_{i}$, and afterwards at each resident change time $s_{j}$, its right slope $h_{i}^{\prime}(\cdot+)$ gets reduced by the slope increment of the resident at this time. This way, $h_{i}(s_{j})$ can be computed inductively for all $s_{j}>t_{i}$; when we first obtain a negative number, we check when precisely between $s_{j-1}$ and $s_{j}$ the trajectory $h_{i}$ hits zero, and from that time on we let $h_{i}$ be identically zero.
@@ Acknowledgements
The authors thank Felix Hermann for interesting discussions and comments. He provided the first construction proving Proposition 9, which was somewhat different from the one appearing in our proof.

### v2

@@ 4 Proofs
Since $\Lambda(\alpha):=\log\mathbb{E}(\mathrm{e}^{\alpha E_{1}})<\infty$ for $\alpha>0$ small (which also follows from the stochastic domination by a geometric random variable), Cramér’s theorem [10, Section 2.2] implies that if $\delta^{\prime}$ is sufficiently small, then
@@
and together with the fact that $\Lambda^{\prime}(0)=m_{1}$ and $\Lambda^{*}(m_{1})=0$, the strict negativity in (4.4) follows from the assertion of [10, Exercise 2.2.24] for $\delta^{\prime}>0$ small enough. We conclude that for all sufficiently large $t$,
@@
as wanted. ∎
Proof of Corollary 3.7.
The analogy between Corollary 3.7 and [16, Theorem 2.8] that we anticipated in Remark 3.8 is the following. $(E_{i})_{i\geq 1}$ is a renewal process (see the proof of Proposition 3.5 for details), just as $(L_{i})_{i\geq 1}$, and
$\widehat{K}_{t}=\sum_{i\geq 0\colon E_{i}\leq t}K_{E_{i}}-K_{E_{i-1}}=\sum_{i\geq 0}K_{E_{i}}\mathds{1}_{[E_{i},E_{i+1})}(t)$
is a renewal reward process with renewal times $E_{i}$, i.e. it has i.i.d. increments at the renewal times $E_{i}$ that are independent of the $E_{i}$’s, analogously to the process
$\widehat{F}(t)=\sum_{i\geq 0\colon L_{i}\leq t}F(L_{i})-F(L_{i-1})=\sum_{i\geq 0}F(L_{i})\mathds{1}_{[L_{i},L_{i+1})}(t)$
defined in [16, Section 5.2], which is a renewal reward process with renewal times $L_{i}$. For $t\geq 0$ let $m(t)$ be such that $E_{m(t)}\leq t<E_{m(t)+1}$. Then we have
$\frac{K_{E_{m(t)}}/m(t)}{E_{m(t)+1}/m(t)}\leq\frac{K_{t}}{t}\leq\frac{K_{E_{m(t)+1}}/m(t)}{E_{m(t)}/m(t)},$
analogously to [16, Equation (5.9)]. Since
  • $m(t)\to\infty$ as $t\to\infty$,
  • $E_{m}$ is a sum of $m$ i.i.d. copies of $E_{1}$ which have finite (and clearly strictly positive) expectation (cf. the proof of Proposition 3.5),
  • and $K_{E_{m}}$ is a sum of $m$ i.i.d. copies of $K_{E_{1}}$,
both the left- and the right-hand side of (4.4) converge almost surely to $\mathbb{E}(K_{E_{1}})/\mathbb{E}(E_{1})$. Hence, in order to finish the proof of the corollary, it suffices to show that $\mathbb{E}(K_{E_{1}})\in(0,\infty)$. The fact that $\mathbb{E}(K_{E_{1}})<\infty$ follows from (4.4) for $i=1$. Finally, $\mathbb{E}(K_{E_{1}})$ is positive because $K_{E_{1}}$ is almost surely at least 2. Indeed, if the input of the PIT is nondegenerate, then between times 0 and $E_{1}$, the slope of the initial resident trajectory must become negative, while the slope of the resident at time $E_{1}$ must become zero. ∎
@@ 5.1 Ancestral relations and fixations
According to [16, Section 6.1], for $i\geq 0$ we say that $h_{i}$ fixes (or reaches fixation) if there exists $t_{0}>0$ such that for all $t>t_{0}$, the trajectory $h_{i}$ is an ancestor (i.e., parent, or parent of parent etc.) of all trajectories $h_{j}$ such that $h_{j}(t)>0$. If the $i$-th trajectory is alone in the system with its parent, $h_{i}$ reaches height 1 at time $t_{i}+\frac{1}{a_{i}}$ and immediately fixes. But when other trajectories are present, due to clonal interference with them, this is not always the case: The slope of the $i$-th trajectory may change due to resident changes after time $t_{i}$.
@@
Given our algorithm in the proof of Theorem 3.3 for the residents, for finite $n$ we can now determine in linear time which trajectories fix. Indeed, the trajectories that fix are precisely the ultimate resident and its ancestors, and this ancestral line can be found in $O(n)$ time using linear search. We can search for the parent of the ultimate resident starting from the end of the list in the output of the algorithm, and then we can search for the parent of this parent continuing backwards etc.
5.2 Relation to Gerrish–Lenski type heuristics
The Gerrish–Lenski heuristics [15] and its refined version [1] provide a heuristic prediction regarding which mutations reach fixation in the population. In the context of the PIT, using these heuristics one can estimate the speed of adaptation explained in Section 1. Here we provide an informal introduction of these heuristics and their purposes, we discuss some issues with their applicability, and we explain how our linear-time algorithm for the residents can help resolve such issues. For a more formal summary of the heuristics, we refer the reader to [16, Section 6.2].
Recall the notion of solitary resident changes from Remark 3.8 and call a trajectory who becomes resident at a solitary resident change time a solitary resident. This notion easily generalizes to any system of interacting trajectories with nondegenerate input. Consider a nondegenerate $\beth=((t_{i},a_{i}))_{1\leq i<n+1}$ with $n\in\mathbb{N}_{0}\cap\{\infty\}$. In this context, the refined Gerrish–Lenski heuristics [1] asserts that for $1\leq i<n+1$, the $i$-th trajectory of the system of interacting trajectories $\mathbb{H}(\beth)$ born at time $t_{i}$ with initial slope $a_{i}$ becomes solitary resident at time $t_{i}+\frac{1}{a_{i}}$ unless
  1. (i)
    there exists some $1\leq j<i$ such that $t_{j}<t_{i}<t_{j}+\frac{1}{a_{j}}$ and $a_{j}\geq a_{i}$, or
  2. (ii)
    there exists some $i<j<n+1$ such that $t_{i}<t_{j}<t_{i}+\frac{1}{a_{i}}$ and $a_{j}>a_{i}$,
and otherwise it never becomes resident. If situation (i) occurs, then we call the $j$-th trajectory a killer of the $i$-th trajectory from the past, while in situation (ii) we call the $j$-th trajectory a killer of $i$-th trajectory from the future. The original Gerrish–Lenski heuristics [15] asserts that the $i$-th trajectory of $\mathbb{H}(\beth)$ becomes solitary resident at time $t_{i}+\frac{1}{a_{i}}$ unless (ii) occurs and otherwise it never becomes resident. The relation between (solitary) resident changes and fixation (cf. Section 5.1) is that only trajectories that become resident can fix and being an ancestor of a solitary resident trajectory is equivalent to eventual fixation. See [16, Section 6.1] for a justification and further details. Note that if all $a_{i}$’s are the same, then there are no killers from the future and every resident change is solitary.
Applying these heuristics sequentially over the birth times, one can guess which trajectories become (solitary) resident. One first asks if the first trajectory becomes resident, if the heuristics says no, one discards the first trajectory and moves on to the second trajectory etc., taking only non-discarded trajectories into account as potential killers from the past. It was implicitly already included in [1] and later also in [16] that if all $A_{i}$’s are equal, then this sequential application of the refined Gerrish–Lenski heuristics provides the correct set of indices of trajectories becoming (solitary) resident. This procedure yields the correct formula for the speed of adaptation originating from the thinning heuristics introduced in [1, Section 3.1], which was described in terms of the PIT in [16, Proposition 2.10]. On the other hand, in the following example we argue that if the $A_{i}$’s are not almost surely constant, then the sequential application of the refined Gerrish–Lenski makes errors almost surely. Our argument is a slight generalization of the example included in [16, Figure 7], illustrated by Figure 1.
Example 5.1.
Assuming that the $A_{i}$’s are not almost surely constant, for any fixed $\varepsilon>0$ there exist some real numbers $a,b$ such that $0<3\varepsilon<a<b<b+\varepsilon$ and $\mathbb{P}(A_{1}\in(a-\varepsilon,a]),\mathbb{P}(A_{2}\in[b,b+\varepsilon)),\mathbb{P}(A_{3}\in(b-\varepsilon,b])$ are all positive. (E.g. the case when $\gamma$ is a discrete measure supported on two points: $a$ and $b$, we choose $A_{1}=a$ and $A_{3}=A_{2}=b$ for any $\varepsilon$.) Now, with positive probability we have $a-\varepsilon<A_{1}\leq a$, $b-\varepsilon<A_{3}\leq b\leq A_{2}<b+\varepsilon$, $T_{1}<T_{2}<T_{1}+\frac{1}{A_{1}}<T_{3}<T_{2}+\frac{1}{A_{2}}<T_{3}+\frac{1}{A_{3}}<T_{4}$, and
$t^{\prime}:=T_{1}+\frac{1}{A_{1}}+\frac{1-(T_{1}+\frac{1}{A_{1}}-T_{2})A_{2}}{A_{2}-A_{1}}>T_{3}+\frac{1}{A_{3}}.$
Then, since $T_{2}+\frac{1}{A_{2}}>T_{1}+\frac{1}{A_{1}}$, the first trajectory becomes resident at time $T_{1}+\frac{1}{A_{1}}$, so that the slope of the second trajectory gets reduced to $A_{2}-A_{1}\leq b+\varepsilon-a+\varepsilon=b-a+2\varepsilon$. Now, we have $0<b-a+2\varepsilon<b-\varepsilon\leq A_{3}$, and since condition (5.1) holds, at time $T_{3}+\frac{1}{A_{3}}$ the third trajectory becomes resident and thus the second one, which would become resident at time $t^{\prime}$ if there were no further kinks after time $T_{1}+\frac{1}{A_{1}}$, never becomes resident. Here, we also used that $T_{4}>T_{3}+\frac{1}{A_{3}}$ and thus the slope of the third trajectory stays equal to $A_{3}$ until the trajectory becomes resident. Thus, the first and the third trajectory fix eventually, while the second does not. Now, since $A_{2}>A_{1}$ and $T_{1}<T_{2}<T_{1}+\frac{1}{A_{1}}$, the second trajectory is a killer of the first one from the future, and since $A_{2}\geq A_{3}$ and $T_{2}<T_{3}+\frac{1}{A_{3}}$, it is also a killer of the third one from the past. On the other hand, the second trajectory has no killer from its past or future (where we also used that $T_{2}+\frac{1}{A_{2}}<T_{4}$). Thus, a sequential application of the refined Gerrish–Lenski heuristics provides the wrong prediction that the second trajectory becomes resident at time $T_{2}+\frac{1}{A_{2}}$, and that the first and the third one never become resident.
[…]
Figure 1: Illustration of Example 5.1 in the case $A_{2}=A_{3}$. The smooth lines represent the true fate of the trajectories of the PIT. The dashed lines show how the zeroth (orange), second (red) and third (green) trajectory would evolve after time $T_{1}+1/A_{1}$ in absence of the first (blue) one. The sequential application of the heuristics suggests this evolution since the first trajectory has a killer in its future and hence it gets discarded. The slope of the dashed green trajectory becomes zero at time $T_{2}+1/A_{2}$ (since $A_{3}=A_{2}$; in the case $A_{3}<A_{2}$ it would turn negative). Without the additional kink at time $T_{3}+1/A_{3}$, the smooth red trajectory would continue as the dotted one, reaching height 1 at time $t^{\prime}$ introduced in (5.1).
Since whether such a configuration of trajectories occurs between two consecutive solitary resident changes is independent of whether it occurs between other consecutive solitary resident changes, the time until the first occurrence of the event can be bounded by a geometrically distributed random variable and it is therefore almost surely finite. We conclude that almost surely, the sequential application of the refined Gerrish–Lenski heuristics leads to errors infinitely often.
Remark 5.2.
Under similar assumptions on the birth times, in [16, Figure 7] it was assumed that $A_{1}=a,A_{2}=b$ and $A_{3}=c$ satisfy $a,c<b$ and $a+c>b$. This leads to similar conclusions but does not cover the case when $\gamma$ is a discrete measure supported on two points (there we need to choose $c=b$). Further, in Example 5.1 we allowed for $\varepsilon$-errors to cover the case when $\gamma$ has less than two atoms.
Based on Example 5.1, for non-constant initial slopes we recommend using our linear-time algorithm instead of the (refined) Gerrish–Lenski heuristics for determining the resident fitness and the sequence of resident trajectories. It yields a correct result and it increases the runtime by at most a constant factor.
We nevertheless believe that other applications of Gerrish–Lenski type heuristics can still be useful. While our algorithm is suitable for providing a Monte Carlo estimate on the speed of adaptation in the PIT (cf. Section 1) and similarly also on the distribution of the resident change times or of the initial fitnesses of residents or fixing trajectories, finding explicit formulas for these quantities still seems difficult for a general $\gamma$. Unlike the original Gerrish–Lenski heuristics, the refined one already seems to provide quite a good numerical approximation for the speed of adaptation (cf. [16, Figure 6]) given in terms of simple explicit integrals (see [16, Section 6.2]), although it tends to underestimate the speed slightly. Further refinements of the heuristics taking into account not only pairwise but also higher-order interactions between trajectories and distinguishing between solitary residency and fixation may improve the precision and help understand better how clonal interference slows down adaptation.
5.3 Computing the height function
Recall that $t\mapsto\varrho^{(\beth)}(t)$ specifies the resident type at time $t\geq 0$, and its array representation can be computed in a runtime of $O(n)$ as in Section 4.2. This algorithm also produces all trajectories in the continued lines representation, each given by a tuple $(t_{i},p_{i},b_{i},y_{i})$ with birth time $t_{i}$, parent index $p_{i}$, slope $b_{i}$, and initial height $y_{i}$. Using the output of this algorithm, the following computations can be made.
The $i$-th ray in the continued lines representation is $l_{i}(t)=y_{i}+b_{i}(t-t_{i})$ for $t\geq t_{i}$. From this, in the original system of interacting trajectories, we have $h_{i}(t)=l_{i}(t)-(l_{p_{i}}(t)-1)$ if this quantity is positive and $h_{i}(t)=0$ otherwise. This can be evaluated at any given $t\geq 0$ in $O(1)$.
@@ 5.3 Computing the height function
Moreover, the entire piecewise linear trajectory $t\mapsto h_{i}(t)$ can be determined in $O(n)$ time as follows: at time $t_{i}$ it starts with slope $a_{i}$, and afterwards at each resident change time $s_{j}$, its right slope $v_{i}(\cdot)$ is reduced by the slope increment of the resident at this time. This way, $h_{i}(s_{j})$ can be computed iteratively from the residency array for all $s_{j}>t_{i}$; when we first obtain a negative number, we check when precisely between $s_{j-1}$ and $s_{j}$ the trajectory $h_{i}$ hits zero, and from that time on we let $h_{i}$ be identically zero.
Finally, to compute $h_{i}(t)$ for all $i$, one may apply the previous procedure independently for each $i$, resulting in an $O(n^{2})$ algorithm. If it is known in advance that each trajectory undergoes at most $O(k)$ kinks, a more efficient approach is to merge the residency array (keyed by residency start time) with the trajectory array (keyed by birth time) and, at each residency change, keep track of and update only the trajectories currently alive. As each trajectory experiences at most $O(k)$ kinks, this yields an algorithm of runtime $O(kn)$.
[…]
For a lower bound, consider an unordered list of pairwise distinct integers $\{x_{1},\dots,x_{n}\}$. For $\beth=\{(x_{1},2),\dots,(x_{n},2)\}$, all trajectories become resident in increasing order of $x_{i}$. Thus, any algorithm that computes the array representation of the piecewise constant functions $t\mapsto\varrho^{(\beth)}(t)$ or $t\mapsto f(t)$ implicitly sorts the $x_{i}$, which in the comparison model requires $\Omega(n\log n)$ time.
Furthermore, the parent of each trajectory $(x_{i},2)$ is the one immediately preceding it in the sorted order of the $x_{i}$. Selecting the maximum $x_{i}$ and repeatedly taking the parent recovers the entire parent–child chain, yielding the sorted sequence in strictly decreasing order of the $x_{i}$. This means that computing the parent-child relationships between the trajectories is as least as hard as sorting. In the comparison model this establishes a lower bound of $\Omega(n\log n)$.
[…]
5.5 Potential extension of the algorithm to the case of moderate selection
Consider a variant of the Moran model where for some $b\in(0,1)$ fixed, the random fitness advantages $A_{i}$ are multiplied by the vanishing factor $N^{-b}$ depending on the total population size $N$. In this setting, mutants typically survive initial fluctuations with probability of order $N^{-b}$ (cf. [3, 4] for more precise statements in this direction). It was conjectured in [16, Section 7.2] based on simulations and heuristic computations that if one still chooses the per generation mutation rate as $\Theta(1/\log N)$ and speeds up time by a factor of $N^{b}\log N$, the system of rescaled logarithmic frequencies of mutants surviving the initial fluctutations converges to a version of the PIT where trajectories start from height $b$ instead of 0 upon birth and become zero after reaching height $b$ from above again (and otherwise they follow the dynamics described in Section 2). We note that the case when fitness advantages are proportional to $N^{-b}$ for some $b\in(0,1)$ is generally referred to as the case of moderate selection, cf. [3, 4].
In the variant of the continued lines representation corresponding to this variant of the PIT, rays are born 1b1-b (instead of $1$) below the current height of the resident ray, and reaching the current resident height minus (1b1-b) again corresponds to becoming 0 again. It is clear that our algorithm for determining the resident fitness remains valid for this variant of the continued lines representation, still with linear runtime and with the height of newborn trajectories being computed accordingly.
@@ Acknowledgements
The authors thank Felix Hermann for interesting discussions and comments. He provided the first construction proving Proposition 3.1, which was somewhat different from the one appearing in our proof. In the context of Remark 4.2, he was the first to observe that a slope of (2k1)-(2k-1) occurs in the case $k=2$. The authors also thank Anton Wakolbinger for historical remarks and comments on the Gerrish–Lenski type heuristics, which we relied on in the discussion of Section 5.2.